A 15 feet ladder is leaning against a wall. The base of the ladder is 3 feet from the wall. How high above the ground is the top of the ladder?
Answer by mananth(16946)

The ladder, ground ,and the wall form a right triangle.
ladder = hypotenuse
wall and distance from base = legs
...
15^2-3^2= H^2
225-9= H^2
216 = H62
H= sqrt 216
Height = 14.7 feet
..
m.ananth@hotmail.ca